# IntentCUA: Learning Intent-level Representations for Skill Abstraction and Multi-Agent Planning in Computer-Use Agents

### v1

@@ 1. Introduction
Rule-based macros and RPA systems enabled early forms of computer-use automation. However, they lack adaptability tripathi2018learning; krosnick2022parammacros when compared to recent GUI agents powered by large language models (LLMs) that can interpret screens and generate actions dynamically. Research on GUI agents has rapidly expanded, spanning web, mobile, and increasingly desktop environments zhang2025largelanguagemodelbrainedgui; sager2025comprehensivesurveyagentscomputer. As highlighted by Tang et al. tang2025surveymllmbasedguiagents, automation across all desktop environments remains particularly challenging due to multi-window operations, OS-level shortcuts and APIs, and the need to adapt to frequent updates and complex software ecosystems. Within such environments, achieving robust long-horizon planning and managing multi-context workflows emerge as central challenges that current systems have yet to overcome.
@@
In end-to-end evaluations, IntentCUA achieves a 74.83% task success rate with a Step Efficiency Ratio (SER) of 0.91, outperforming both RL-based (UI-TARS-1.5) and trajectory-centric (UFO2) baselines in success rate, efficiency, and latency. Ablation studies confirm that multi-view intent abstraction and shared plan memory jointly improve execution stability, with the cooperative multi-agent loop providing the largest gains on long-horizon tasks. These results indicate that system-level intent abstraction and memory-grounded coordination are central to reliable desktop automation.
@@ 2.1. Desktop and GUI Automation Agents
Research on GUI automation spans web, mobile, and desktop domains. Web-based agents demonstrate end-to-end navigation under structured DOM feedback but remain restricted to browser environments zhou2023webarena; he2024webvoyager. Mobile and sandboxed agents such as SUGILITE li2017sugilite and AppAgent zhang2025appagent leverage structured knowledge bases or multi-modal reasoning for mobile automation li2020mapping; yang2023mm. Large-scale datasets like AITW enable pretraining of device-level agents rawles2023android.
Desktop environments, however, introduce additional challenges such as multi-application workflows, window switching, and device-level resources. Frameworks including UI-TARS qin2025ui, UFO zhang2024ufo, ScreenAgent niu2024screenagent, and AssistGUI gao2023assistgui extend vision-language models to native GUIs and often rely on planner-critic loops. Despite this progress, these systems are typically evaluated in constrained or schema-aware setups, limiting generalization to real operating systems.
Perception bottlenecks have been addressed by GUI grounding methods, which convert pixels into actionable elements and align them with semantics lu2024omniparser; hong2024cogagent; jiang2025iluvui. While improved perception reduces local errors, it does not prevent agents from repeatedly re-solving already completed subroutines when long-horizon intent coherence is lost.
@@ 2.2. Agents Leveraging Interaction Traces
Interaction traces encode reusable procedural knowledge. Macro-mining clusters demonstrations into recurrent procedures huang2024automatic, while process-mining techniques transform logs into labeled schemas for interoperability fani2023llms; choi2022enabling. Large-scale corpora such as OS-ATLAS support pretraining of perception models over millions of GUI elements wu2025osatlas.
Offline reinforcement learning has also been explored for device agents, primarily focusing on low-level control policies song2023navigating. Recent systems such as AppAgentv2 li2025appagentv2advancedagent introduce hierarchical feedback reinforcement learning for mobile UIs, and AgentBank song2024agentbank demonstrates that large-scale trajectory-tuning across tens of thousands of demonstrations can improve the generalization of LLM agents. UI-TARS-1.5 ui-tars-15-seed further exemplifies this line of work, employing reinforcement learning to improve self-evolving GUI automation.
These works confirm the utility of traces for replay, pretraining, and hierarchical control. Yet most approaches stop short of compressing trajectories into structured intent abstractions directly usable by planners. Replay-based reuse or large-scale tuning tends to improve policy robustness but often reintroduces redundancy and amplifies error accumulation in long-horizon execution redis2024skill. Their reliance on controlled environments further limits applicability to broad computer use tasks, where rewards are difficult to define and action consequences cannot easily be suppressed sager2025comprehensivesurveyagentscomputer.
@@ 2.3. Plan Memory, Intent Identification, and Skill Abstraction
Memory mechanisms enhance robustness by enabling reuse of prior experience. Reflexion shinn2024reflexion, Conversational Memory wang2023conversational, and Contextual Experience Replay cer2025 retrieve past trajectories, manuals, or reflections to guide future decision-making cai2023low. These methods primarily emphasize replay rather than skill abstraction. SkillAct liu2024skillact, for example, shows that prompting with skill abstractions can improve interactive performance, but it lacks structured intent prototypes or hierarchical plan memory. Similarly, UFO2 zhang2025ufo2 manages app-specific documents and user demonstrations as trajectories, yet does not elevate them into higher-level intent representations.
Parallel work explores identifying user goals from UI logs rebmann2024recognizing; li2020mapping and learning representation-based embeddings of screens (e.g., Screen2Vec li2021screen2vec, Aria-UI yang2024aria). GUI grounding techniques lu2024omniparser; li2024ferret indirectly support intent recognition by reducing perceptual errors. More recent efforts such as GUI-Explorer xie2025guiexplorer and Dynamic Planning for GUI Automation zhang2024dynamicplanningGUI emphasize schema-less exploration or adaptive plan decomposition, while dependency modeling from user demonstrations provides cognitive structures useful for task automation yin2025cognitivedependencies. However, these methods rarely induce modular skills transferable across heterogeneous workflows.
Our approach complements these directions by introducing a multi-view latent representation that integrates environment, action, and description signals. Skills are represented as intent prototypes stored hierarchically in plan memory and retrieved at runtime to augment partial plans. This design supports robust long-horizon execution and improves cross-application generalization, complementing advances in both trace mining and agent memory song2024visiontasker; gao2023assistgui.
@@ 4.2. Optimizing steps by memory & feedback loop
The InjectHint function searches the plan memory for previous plan units whose representations $z^{(pu)}$ are most similar to the current one, and uses their traces as contextual hints guiding the next execution steps.
@@ 5.1. Evaluation Setup
For model training, we used a learning rate of $1\times 10^{-3}$, with $\lambda_{\mathrm{pred}}=0.1$, $\lambda_{\mathrm{rec}}=0.05$, and a contrastive temperature $\tau=0.1$. All agents use the same atomic GUI action interface and identical timeout policies. We report task success (74.83%), average completion ratio (91.14%), Step Efficiency Ratio(successful steps / actual execution steps; higher is better). Differences are reported in percentage points (pp).
@@ 5.2. End-to-End Execution of Ablated Models
We ablate five variants to quantify the contribution of multi-view $IG/SG$ clustering, plan memory reuse, and $SG$ derived skill hints. Long-horizon tasks are those with $\geq$25 low-level actions.
Table 1. Ablation of components. We report task success (%) and completion (%) for: no traces; Greedy retrieval; intent-level rep; + plan-memory (reuse $G$); + SG-based skill hints, and full system without intent-level representation.
@@
No User Traces 22.73 33.78
Greedy retrieval 46.43 57.41
intent-level rep 54.64 77.56
IntentCUA w/o intent-level rep 53.85 81.23
IntentCUA w/o Skill Hint ($S_{SG}$) 62.51 85.00
IntentCUA(Full System) 74.83 91.14
For Table 1, Completion is the averaged plan completion: for each task, we compute it (executed low-level steps/steps in the synthesized plan) and then average over tasks. This measures how far execution progresses along the agent’s own plan, complementing binary success. Ablated model, greedy retrieval picks the top-3 most similar past intent units without filtering. This helps inject traces but often pulls in mismatched routines, causing drift in long sequences. $IG/SG$ gating restricts retrieval to representation-learned intent groups and subgroups, ensuring retrieved traces align with the current intent rather than arbitrary matches.
As shown in Table 1, starting from a no-trace baseline (22.73%), simply adding user traces with greedy retrieval lifts success by +23.7 percentage points (pp) and increases plan completion +23.65 pp. The gain stems from injecting user-specific web/app workflows, which improves both screen interpretation and execution-flow understanding, though long sequences still drift. Adding $IG/SG$ gating confines retrieval to intent-coherent groups/subgroups, yielding more accurate partial plans and increasing success +8.21 pp over greedy while markedly boosting completion +20.13 pp. Plan-memory reuse further raises success +7.87 pp and completion +7.44 pp by binding cached stored plans $G$ rather than re-synthesizing. Finally, skill hints fill missing plan units/steps on partial matches, improving success to 74.83% (+12.32 pp) and completion to 91.14% (+6.14 pp).
@@ 5.3. Case Study
To illustrate the behavior of the proposed framework, we present a case where the user asks the system to summarize a previously viewed video about lightweight-ML and record the result in a personal workspace. The system interprets this request as a multi-intent task involving retrieval, reasoning, and organization across different applications.
@@
Figure 4 shows how the system handles this command through three main processes. First, the Planner interprets the command and retrieves relevant traces from the plan memory that include traces involving a video platform, AI chat-bot, and workspace app and, crucially, recovers the video’s source from the user’s prior viewing trace. Based on these retrieved traces, it constructs a structured breakdown into four plan units, reflecting the GUI context, the abstracted action sequence, and the personalized information required to fulfill the corresponding intent. Each unit is grounded by the Plan-Optimizer into executable GUI actions, while the Critic monitors progress and triggers local re-planning when necessary. Through this hierarchical reasoning and memory-guided skill retrieval, the system autonomously executes a multi-application task, demonstrating consistent intent understanding across heterogeneous software environments.
In contrast, when a pop-up window unexpectedly appears during execution, the system encounters a failure case. Because the underlying GUI components remain hidden beneath the pop-up, the grounding module fails to detect them, leading the Critic to assume that the intended component no longer exists on the screen. This case highlights a limitation of script-based GUI grounding, which lacks visual awareness of interface changes such as transient pop-ups or overlays.
@@ 6.1. Robust Long-Horizon Planning Efficiency
As shown in Figure 5, IntentCUA maintains stable performance even as task length grows: 85.9% at 10–15 steps, 72.5% at 15–20, and 65.0% at 20–25, while still retaining 42.9% beyond 30 steps. Both baselines, in contrast, decline sharply after 20 steps, dropping below 20This gradual degradation indicates that IntentCUA’s planning remains consistent and resistant to drift even in extended workflows spanning multiple windows and applications.
@@ 7. Conclusion
While IntentCUA shows consistent reasoning and cross-application generalization, several aspects remain open for refinement. Retrieval efficiency may fluctuate as the plan memory grows, though this mainly affects latency rather than accuracy. Graph-based retrieval and lightweight vision cues could further enhance robustness, allowing the system to adapt more smoothly to dynamic and visually changing interfaces.
@@ Appendix C Ablation on the Representation Loss
To address concerns about the representation objective, we compare four variants: InfoNCE-only, InfoNCE + cross-view prediction, InfoNCE + reconstruction, and the full loss. We report (1) size-weighted density separation (inter/intra) over all HDBSCAN subgroups in Table 3 and (2) semantic purity measured as majority ENV/ACT ratios within each subgroup in Table 4.

### v2

# arXiv:2602.17049v2[cs.AI] 27 Feb 2026

@@ 1. Introduction
Rule-based macros and RPA systems enabled early forms of computer-use automation. However, they lack adaptability tripathi2018learning; krosnick2022parammacros when compared to recent GUI agents powered by large language models (LLMs) that can interpret screens and generate actions dynamically.
Research on GUI agents has rapidly expanded, spanning web, mobile, and increasingly desktop environments zhang2025largelanguagemodelbrainedgui; sager2025comprehensivesurveyagentscomputer. As highlighted by Tang et al. tang2025surveymllmbasedguiagents, automation across all desktop environments remains particularly challenging due to multi-window operations, OS-level shortcuts and APIs, and the need to adapt to frequent updates and complex software ecosystems. Within such environments, achieving robust long-horizon planning and managing multi-context workflows emerge as central challenges that current systems have yet to overcome.
@@
In end-to-end evaluations, IntentCUA achieves a 74.83% task success rate with a Step Efficiency Ratio (SER) of 0.91, outperforming both RL-based (UI-TARS-1.5 ui-tars-15-seed) and trajectory-centric (UFO2 zhang2025ufo2) baselines in success rate, efficiency, and latency. Ablation studies confirm that multi-view intent abstraction and shared plan memory jointly improve execution stability, with the cooperative multi-agent loop providing the largest gains on long-horizon tasks. These results indicate that system-level intent abstraction and memory-grounded coordination are central to reliable desktop automation.
@@ 2.1. Desktop and GUI Automation Agents
GUI automation spans web, mobile, and desktop domains. Web agents such as WebArena and WebVoyager operate under structured DOM feedback zhou2023webarena; he2024webvoyager, but real desktop environments lack such schema-level constraints and require cross-application coordination.
Desktop benchmarks like OSWorld xie2024osworldbenchmarkingmultimodalagents tasks are typically long-horizon, requiring stable execution 10–20 sequential steps. This makes execution latency-sensitive and error-prone. As step count increases, local perception errors compound and agents often enter loops of repeated or failed actions.
Recent desktop agents such as UI-TARS qin2025ui, UFO zhang2024ufo, ScreenAgent niu2024screenagent extend vision-language models with planner–critic loops. However, surveys highlight a persistent challenge: determining actions that align with specific user contexts and preferences in dynamic, interruption-prone interfaces tang2025surveymllmbasedguiagents. Even with improved GUI grounding lu2024omniparser; hong2024cogagent; jiang2025iluvui, intent drift and redundant re-planning remain common in long-horizon workflows.
These findings indicate that stable long-horizon planning, rather than perception alone, remains the key bottleneck for reliable desktop automation.
@@ 2.2. Agents Leveraging Interaction Traces
One approach to addressing long-horizon instability is to learn directly from large-scale interaction traces.
Macro-mining and process-mining techniques cluster demonstrations into recurrent procedures or labeled schemas huang2024automatic; fani2023llms; choi2022enabling. Large-scale corpora such as OS-ATLAS support perception pretraining across millions of GUI elements wu2025osatlas. Offline reinforcement learning has also been explored for device agents song2023navigating, while systems such as AppAgentv2 li2025appagentv2advancedagent, AgentBank song2024agentbank, and UI-TARS-1.5 ui-tars-15-seed leverage hierarchical feedback or large-scale trajectory tuning to improve control robustness.
These works demonstrate that interaction traces improve policy generalization and low-level stability. However, most approaches operate at the trajectory or action level, emphasizing replay or large-scale tuning rather than structured intent abstraction. As a result, redundancy and error accumulation often persist in long-horizon execution redis2024skill, and reliance on controlled environments or explicit reward signals limits applicability to open-ended desktop workflows sager2025comprehensivesurveyagentscomputer.
@@ 2.3. Plan Memory, Intent Identification, and Skill Abstraction
A complementary direction enhances robustness through memory retrieval and skill abstraction.
Memory-based methods such as Reflexion shinn2024reflexion, Conversational Memory wang2023conversational, and Contextual Experience Replay cer2025 retrieve prior trajectories, manuals, or reflections to guide future decisions cai2023low. Skill-level prompting approaches such as SkillAct liu2024skillact show that abstracted routines can improve interactive performance, while UFO2 zhang2025ufo2 manages app-specific demonstrations as reusable references.
Parallel work investigates intent recognition from UI logs rebmann2024recognizing; li2020mapping and representation learning of screens (e.g., Screen2Vec li2021screen2vec, Aria-UI yang2024aria), while GUI grounding methods reduce perceptual ambiguity lu2024omniparser; li2024ferret. More recent systems explore adaptive planning and dependency modeling from demonstrations zhang2024dynamicplanningGUI; yin2025cognitivedependencies.
Despite these advances, structured and hierarchical skill abstractions that remain transferable across heterogeneous desktop workflows are still relatively underexplored. As a result, maintaining stable long-horizon execution under dynamic user contexts continues to be an active area of research. Our approach complements these directions by learning multi-view intent representations that integrate environment, action, and description signals. Skills are stored as hierarchical intent prototypes in plan memory and retrieved to augment partial plans, supporting stable long-horizon execution song2024visiontasker; gao2023assistgui.
@@ 4.2. Optimizing steps by memory & feedback loop
The InjectHint function searches the plan memory for previous plan units whose representations $z^{(pu)}$ are most similar to the current one, and uses their traces as contextual hints guiding the next execution steps. Example of Planner-Plan-Optimizer-Critic interactions is in Appendix B.
@@ 5.1. Evaluation Setup
All agents use the same atomic GUI action interface and identical timeout policies. We report task success (74.83%), average completion ratio (91.14%), Step Efficiency Ratio(successful steps / actual execution steps; higher is better). Differences are reported in percentage points (pp).
@@ 5.2. End-to-End Execution of Ablated Models
We ablate each component to analyze its impact on long-horizon planning stability and execution depth.
Table 1. Component-wise ablation on planning. We report task success (%) and averaged plan completion (%).
@@
$B$ 22.73 33.78
$B+T_{g}$ 46.43 57.41
$B+T_{SG}+Z$ 54.64 77.56
$B+T_{SG}+S_{SG}+PM$ 53.85 81.23
$B+T_{SG}+Z+PM$ 62.51 85.00
$B+T_{SG}+Z+S_{SG}+PM$ 74.83 91.14
For Table 1, Completion denotes averaged plan completion (executed steps / synthesized plan steps), measuring execution progress beyond binary success.
We denote components as: $B$ (baseline planner–executor), $T_{g}$, $T_{SG}$ (greedy vs. $IG/SG$-gated trace retrieval), $Z$ (intent-level representation), $S_{SG}$ (subgroup skill hints), $PM$ (plan memory reuse)
[…]
@@ 5.3. Case Study
To illustrate the framework, we consider a task where the user asks the system to summarize a previously viewed lightweight-ML video and record the result in a personal workspace. This requires retrieval, reasoning, and coordination across multiple applications.
@@
Figure 4 shows the execution process. The Planner retrieves relevant traces from plan memory, including prior interactions with the video platform, AI chatbot, and workspace application, and reconstructs the video source from the user’s history. It decomposes the request into structured intent-level plan units, each grounded into executable GUI actions by the Plan-Optimizer, while the Critic monitors progress and handles local inconsistencies. Through hierarchical reasoning and memory-guided skill retrieval, the system completes the multi-application task while maintaining intent coherence.
A failure case occurs when an unexpected pop-up appears during execution. Because underlying components become occluded, the grounding module fails to detect them, leading to incorrect retries. This highlights a limitation of script-based GUI grounding under transient interface changes.
@@ 6.1. Robust Long-Horizon Planning Efficiency
As shown in Figure 5, IntentCUA maintains stable performance even as task length grows: 85.9% at 10–15 steps, 72.5% at 15–20, and 65.0% at 20–25, while still retaining 42.9% beyond 30 steps. Both baselines, in contrast, decline sharply after 20 steps, dropping below 20%. This gradual degradation indicates that IntentCUA’s planning remains consistent and resistant to drift even in extended workflows spanning multiple windows and applications.
@@ Appendix A Encoder Details
For model training, we used a learning rate of $1\times 10^{-3}$, with $\lambda_{\mathrm{pred}}=0.1$, $\lambda_{\mathrm{rec}}=0.05$, and a contrastive temperature $\tau=0.1$.
@@ Appendix C Ablation on the Representation Loss
To address concerns about the representation objective, we compare four variants: InfoNCE-only, InfoNCE + cross-view prediction, InfoNCE + reconstruction, and the full loss.
$\mathrm{Sep}_{\mathrm{w}}=\frac{\sum\limits_{IG}\sum\limits_{SG\in IG}|SG|\;\frac{\mathrm{Inter}(SG)}{\mathrm{Intra}(SG)}}{\sum\limits_{IG}\sum\limits_{SG\in IG}|SG|}$ (7)
$\mathrm{Intra}(SG)$ $=\frac{1}{|SG|}\sum_{x\in SG}d(x,\mathbf{c}_{SG})$ (8)
$\mathrm{Inter}(SG)$ $=\min_{SG^{\prime}\neq SG}d(\mathbf{c}_{SG},\mathbf{c}_{SG^{\prime}})$
$\mu^{X}$ $=\frac{1}{|\mathcal{SG}|}\sum_{IG}\sum_{SG\in IG}\mathrm{Purity}^{X}(SG)$ (9)
$\sigma^{X}$ $=\sqrt{\frac{1}{|\mathcal{SG}|}\sum_{IG}\sum_{SG\in IG}\left(\mathrm{Purity}^{X}(SG)-\mu^{X}\right)^{2}}$
$\mathrm{Purity}^{X}(SG)=\frac{\max_{c}\left|\{\,t\in SG:\mathrm{tag}^{X}(t)=c\}\right|}{|SG|}$ (10)
We report (1) size-weighted density separation (inter/intra: Eq 8) in Eq 7 over all HDBSCAN subgroups in Table 3 and (2) semantic purity(Eq 9) measured as majority ENV/ACT ratios(Eq 10) within each subgroup in Table 4.